OpenStudy (anonymous):

convert the angle 11pie over to degree

OpenStudy (acxbox22):

2pi equals 360 degrees

OpenStudy (anonymous):

\[11\pi?\] from radians to degrees? pi radians = 180 degrees.

OpenStudy (anonymous):

Welcome to OpenStudy! Unfortunately, there is no such thing as a `pie` in mathematics. We do however have a \(\pi\). And it is written as `pi` not `pie`. `pi` means it is in radians unless the teacher was really sneaky and tried to put a degree sign with the radian. As mentioned above, \(\sf\pi=180 ~degrees\) However, if you multiply both sides by 2, so that way it is still in a ratio, then: \(\sf 2\pi = 360~degrees\) Also do note that, 360 degrees will make one whole revolution.

OpenStudy (matt6288):

if your trying to convert radians to degrees you multiply the radiant value by 180/pi for example 4pi/2 to find the degree you multiply it to 180/pi and you get 360. To me this is the easiest way of converting it. to convert degrees to radiants you flip it and multiply it to pi/180. Anyway I hope this helped.